OpenStudy (anonymous):

solve by the substitution method 2m+n=5 m-9n=12???

OpenStudy (anonymous):

n=5-2m plug "5-2m" into the 2nd equation instead of "n" (substitution): m-9(5-2m)=12 find m from above, then find n from the first equation (using the m you found).

OpenStudy (anonymous):

does that make sense nikkiem1 ?

OpenStudy (anonymous):

2m+n=5 n=5-2m m-9n=12 m-9(5-2m)=12 m-45+18m=12 m+18m=57 19m=57 m=3

OpenStudy (anonymous):

do i have to collect like terms?

OpenStudy (anonymous):

as you see above in Kush's steps - collection of like terms is being done.

OpenStudy (anonymous):

For example: these steps are exactly that - collection of like terms: m-45+18m=12 m+18m=57 19m=57

OpenStudy (anonymous):

now that you know that m=3 can you find n=?